OpenStudy (anonymous):

A street light is mounted at the top of a 15 foot-tall pole. A man 6ft tall walks away from the pole with the speed of 5ft/s along a straight path. How fast is the tip of his shadow moving when he is 40ft from the pole?

OpenStudy (anonymous):

..........................| ..........................| ...........|..............| ...........|..............| 15 ' T-------|------------ man pole Let x be the distance between the man and the pole. dx/dt=5 Let y be the distance between the man and the tip of his shadow (length of the shadow) Let x+y = z (the distance between the tip of the shadow and the pole) By using similar triangles, z/15 =x / (15-6) z/15 = x/9 z=(15/9) x z=(5/3)x dz/dt = (5/3) dx/dt dz/dt = (5/3)(5)=25/3=8 1/3 Regardless of how far he is from the pole, the tip of his shadow is moving at a rate of 8 1/3 ft /s

OpenStudy (anonymous):

Hope this was helpful, give me medal.

OpenStudy (anonymous):

i don't get it

OpenStudy (phi):



OpenStudy (phi):

I think that is how blast set it up, using similar triangles.

OpenStudy (phi):

if you get that far, the calculus is pretty straight-forward. 15 x = 9 s 15 dx/dt =9 ds/dt or ds/dt = 5/3 dx/dt

OpenStudy (phi):

I am not sure how you figure this out, but the top triangle with base leg x is moving at 5 ft/sec. Meanwhile, the base leg s is moving at ds/dt (it is growing faster than 5 ft/sec because the shadow is growing longer...)